OpenStudy (anonymous):

the vertex of 1/4 over 2(16)

OpenStudy (saifoo.khan):

\[\huge \frac{\frac 14 }{2(16)}\]

OpenStudy (saifoo.khan):

that?

OpenStudy (anonymous):

yeah but doesn't it equal to .008 that is to small right.

OpenStudy (anonymous):

??

OpenStudy (saifoo.khan):

No, that doesnt equal to that.

OpenStudy (anonymous):

does .25/32 = .007

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

\[7.8125 \times 10^{-3}\]

OpenStudy (anonymous):

well the x still wont be that small

OpenStudy (saifoo.khan):

there might be some mistake in the equation.

OpenStudy (anonymous):

xthe equation is 16-1/4^2

OpenStudy (saifoo.khan):

where is x?

OpenStudy (anonymous):

16-1/4X^2

OpenStudy (saifoo.khan):

that will be a curve

OpenStudy (anonymous):

I still don;t get it